# Pun Unintended: LLMs and the Illusion of Humor Understanding

### v1

@@ 1 Introduction
Motivated by these limitations, we introduce two new collections of annotated short texts, PunnyPattern and PunBreak, designed specifically to evaluate the robustness of LLMs in pun detection (see Fig. 1). These collections move beyond simple detection, using targeted substitutions and common language patterns to probe whether models can accurately recognize puns’ context and structure, or if they rely on memorization and superficial cues. We publicly release both datasets. [^1] https://github.com/alezanga/punintended
@@ 2 Related Work
Building upon this foundation, the JOKER workshop Ermakova et al. (2023b) introduced multilingual datasets for pun detection, location, interpretation, and translation, encompassing English, French, and Spanish. The JOKER corpus Ermakova et al. (2023a) aimed to address earlier criticisms of imbalance and reliance on superficial cues by providing 3,506 annotated short jokes in English and French, alongside 1,700 negative examples generated via word substitutions. Despite utilizing advanced LLMs such as GPT-3 Brown et al. (2020) and BLOOM Workshop et al. (2023), a custom-trained T5 model Raffel et al. (2019) consistently outperformed others on all tasks, albeit with limited success (detection F1 < 0.6). Other studies have further explored non-English puns Gameiro et al. (2025); Chen et al. (2023); Dsilva (2024).
Complementing these efforts, Sun et al. (2022a) investigated pun explanation by enhancing training data with rationales generated by T5 models. They measured explanation quality using simulatability, defined as the accuracy difference when explanations are provided as additional input vs. when they are not. Their findings emphasized that high-quality human-annotated rationales could improve model performance, although automatically generating such explanations using LMs is challenging. In a related work, Sun et al. (2022b) focus on pun generation by retrieving context and pun words. They propose a system that first retrieves a suitable pair of words, along with their senses, from given contextual words, and then employs a T5 model to generate a pun. However, the success rate of the generated puns was reported to be significantly lower than that achieved by humans.
@@ 4.2 Results
In addition to the comparison LLMs, we also include a RoBERTa-large encoder model Liu et al. (2019) that has been fine-tuned on the PunEval training split. The significant performance difference observed between PunEval (92.5) and the other datasets (64.0 in NAP and 65.7 in JOKER) prompted an investigation to understand the reasons behind this disparity. We analyzed the embedding distribution of the RoBERTa model using t-SNE van der Maaten and Hinton (2008) to visualize the embeddings as points in a 2D space (Fig. 2, see more details in Appendix §A). The visualizations reveal poor separation between subtly altered positive and negative examples in the NAP and JOKER datasets (Fig. 2, right). In contrast, PunEval embeddings show better separation even before fine-tuning (Fig. 2, center), with clusters often corresponding to recurring templates. For instance, the top-left cluster from the middle image represents puns structured as “Old […] never die, they just”, a common pattern for creating puns (e.g., “Old bankers never die, they just lose interest”). This suggests that dataset artifacts or structural similarities in PunEval may facilitate shortcut learning.
@@ 5.2.2 Results
The binary accuracy of each LLM for every substitution subset, using each model’s best prompt from the RQ1 experiments, is shown in Fig. 3. The “pun” subset contains exactly 200 puns (100 het-puns and 100 hom-puns); every other subset contains exactly 200 non-puns representing the various substitutions. Results from single prompts are reported in Appendix §C.4. The “rand sent” results are obtained from a control of 100 randomly generated non-puns to detect any unseen bias toward the pun class; scores exceed 0.8, indicating no such bias in the random sentences.
@@
All models exhibit a bias toward labeling altered examples as puns, reflected in low accuracy across all substitution categories, which contain only negative examples (non-puns). In particular, GPT-4o, the best performer overall, reaches a low of 0.33 on the homophone subset and a high of 0.59 on alt syn. Conversely, all models achieve accuracy of around 0.8 or higher on the negative examples from PunEval and the control group (rand sent) (see Table 11 for details). The homophone subset is the most challenging overall, indicating that phonetic similarity between $w_{p}$ and its replacement degrades the LLMs’ ability to distinguish them in context. This may also be influenced by orthographic similarity, as similar phonetics often imply similar orthography (e.g., stock/stork or dragon/brag gone in Example 3). Pun-word synonym substitutions (pun syn) also generally challenge the models, suggesting they may overlook the effect of the replaced word when it is semantically similar to a more fitting alternative. This suggests that LLMs might be actively “fixing” the pun in their reasoning process, thereby rectifying the alteration. The random and alt syn sets show slightly higher accuracy overall, further indicating that semantic or phonetic similarity to valid puns misleads predictions more than loosely related substitutions. These results suggest that LLMs fail to recognize contexts that should not trigger the pun effect in cases where the target sentence has a typical pun structure.
@@ Results.
Table 3 reports the average PPA over three runs, representing the mean number of words correctly identified in generated rationales for pun pairs on a 0–2 scale. GPT-4o is the top performer, with $\sim$1.5 out of 2 correctly predicted words per pun across all three datasets. It is followed by Llama3.3 and DeepSeek-R1, whose performance is noticeably lower. Mistral has the lowest scores, trailing the next-lowest model, Gemini2, by $\sim$33%. Among the 70B models, Llama3.3 performs best on average. We also observed no significant difference between the two prompts tested. Consistent with our earlier results, scores on the PunEval datasets are generally higher than on JOKER and NAP, where performance drops by approximately 20% and peaks at 1.5 out of 2. Furthermore, a PPA evaluation restricted to true positives (see Appendix Table 12) shows that most models, except Qwen and Mistral, can effectively justify their correct answers. In this specific evaluation, GPT-4o and DeepSeek-R1 emerge as the top performers by a wide margin. These findings motivate our investigation of the main error types in model reasoning, which is addressed in the next subsection.
@@ 6.3 Discussion
Our findings align with prior work showing that LLMs struggle with humor, particularly when it depends on nuanced context or roleplay Quan et al. (2025); Mirowski et al. (2024). These studies note that safety constraints (e.g., the “Harmless, Helpful, Honest” criteria) can cause LMs to misinterpret relational context in potentially offensive situations. These safety mechanisms bias models toward a safer, “washed-out” form of humor, stripping away the surprise and edge on which many jokes depend. Puns, in contrast, are rich in surprise, subtlety, and timing; they are also deeply rooted in human emotions and experiences, qualities that current LLMs cannot fully tap into.
@@
Finally, the frequent pun pair errors point to weaknesses in assessing phonetic and orthographic similarity. These issues are likely tied to tokenization, which can mask morphological elements critical to wordplay, and to limited phonological modeling Liao and Shi (2025). The composition of pre-training corpora and data-cleaning practices may also deprive models of the informal, creative language where puns are common. Because developers rarely disclose detailed preprocessing and corpus information, it is difficult to determine how these practices affect a model’s fluency with such language.
Improving an AI’s grasp of puns and humor in general will likely require deeper human-machine collaboration and greater social awareness. Prior work suggests that human alignment of LMs should shift from a single global standard to a more granular, audience-aware approach Quan et al. (2025); Mirowski et al. (2024). This would help models to adapt to specific audiences and contexts, adjusting appropriateness and fairness criteria accordingly.
@@ Acknowledgments
We thank Felix Chadwick-Smith, Fayeja Farhana, Dhruvil Raval, Kayleigh Shier, and Hannah Trotman for their work as annotators on the manual error analysis of rationales. We also thank the entire Cardiff NLP group for guidance on the project and for feedback on an early draft of the paper.
@@ C.3 RQ2: Pun Language Patterns
Table 10 compares the results obtained using each prompt for PunEval and PunnyPattern. Note that the best prompts (in bold) are the same as those that performed best in our RQ1 (Table 1). Figure 11 visualizes the performance differences for each set of expressions containing language patterns. We interpret bias as being present in models with very high recall but low precision, indicating they tend to classify all instances containing the pattern as puns.
We observe a sharp imbalance of low precision and high recall for the never_die, used, and doctor patterns (Fig. 11), indicating that these patterns tend to mislead LLMs into identifying puns. In contrast, the tom and when patterns do not appear to consistently affect them. While all LLMs tend to produce many false positives with the used, never die, and doctor patterns, results for the other patterns are more varied. For instance, GPT-4o, Qwen2.5, and Mistral3 are more resilient to the daughter pattern compared to the other models. Interestingly, examples containing the word “Tom” produce both false positives and negatives, suggesting that this pattern is unlikely to serve as a reliable shortcut for classifying puns. In contrast, sentences beginning with “When the” are classified more reliably by all models.
@@ C.5 RQ3
This targeted analysis reveals that most models can generate reasonably accurate rationales for the puns they successfully identify. Mistral and Qwen achieve lower scores in this setting, while Gemini’s performance is comparable to Llama’s.

### v2

@@ 1 Introduction
Motivated by these limitations, we introduce two new collections of annotated short texts, PunnyPattern and PunBreak, designed specifically to evaluate the robustness of LLMs in pun detection (see Fig. 1). These collections move beyond simple detection, using targeted substitutions and common language patterns to probe whether models can accurately recognize puns’ context and structure, or if they rely on memorization and superficial cues. We publicly release both datasets [^1] https://github.com/alezanga/punintended.
@@ 2 Related Work
Building upon this foundation, the JOKER workshop Ermakova et al. (2023b) introduced multilingual datasets for pun detection, location, interpretation, and translation, encompassing English, French, and Spanish. The JOKER corpus Ermakova et al. (2023a) aimed to address earlier criticisms of imbalance and reliance on superficial cues by providing 3,506 annotated short jokes in English and French, alongside 1,700 negative examples generated via word substitutions. Despite utilizing advanced LLMs such as GPT-3 Brown et al. (2020) and BLOOM Workshop et al. (2023), a custom-trained T5 model Raffel et al. (2020) consistently outperformed others on all tasks, albeit with limited success (detection F1 < 0.6). Other studies have further explored non‑English puns Gameiro et al. (2025); Chen et al. (2023); Dsilva (2024).
Complementing these efforts, Sun et al. (2022a) investigated pun explanation by enhancing training data with rationales generated by T5 models. They measured explanation quality using simulatability, defined as the accuracy difference when explanations are provided as additional input vs. when they are not. They found that high-quality human-annotated rationales can improve model performance, but automatically generating such explanations with language models remains challenging. In a related work, Sun et al. (2022b) focus on pun generation by retrieving context and pun words. They propose a system that first retrieves a suitable pair of words, along with their senses, from given contextual words, and then employs a T5 model to generate a pun. However, the success rate of the generated puns was reported to be significantly lower than that achieved by humans.
@@ 4.2 Results
In addition to the comparison LLMs, we also include a RoBERTa-large encoder model Liu et al. (2019) that has been fine-tuned on the PunEval training split. The significant performance difference observed between PunEval (92.5) and the other datasets (64.0 in NAP and 65.7 in JOKER) prompted an investigation to understand the reasons behind this disparity. We analyzed the embedding distribution of the RoBERTa model using t-SNE van der Maaten and Hinton (2008) to visualize the embeddings in a 2D space (Fig. 2, see details in Appendix §A). The visualizations reveal poor separation between subtly altered positive and negative examples in the NAP and JOKER datasets (Fig. 2, right). In contrast, PunEval embeddings show better separation even before fine-tuning (Fig. 2, center), with clusters often corresponding to recurring templates. For instance, the top-left cluster from the middle image represents puns structured as “Old […] never die, they just”, a common pattern for creating puns (e.g., “Old bankers never die, they just lose interest”). This suggests that dataset artifacts or structural similarities in PunEval may facilitate shortcut learning.
@@ 5.2.2 Results
The binary accuracy of each LLM for every substitution subset, using each model’s best prompt from the RQ1 experiments, is shown in Fig. 3. The “pun” subset contains exactly 200 puns (100 het-puns and 100 hom-puns); every other subset contains exactly 200 non‑puns representing the various substitutions. Results from single prompts are reported in Appendix §C.4. The “rand sent” results are obtained from a control of 100 randomly generated non-puns to detect any unseen bias toward the pun class; scores exceed 0.8, indicating no such bias in the random sentences.
@@
All models exhibit a bias toward labeling altered examples as puns, reflected in low accuracy across all substitution categories, which contain only negative examples (non-puns). In particular, GPT‑4o, the best performer overall, reaches a low of 0.33 on the homophone subset and a high of 0.59 on alt syn. Conversely, all models achieve accuracy of around 0.8 or higher on the negative examples from PunEval and the control group (rand sent) (see Table 11 for details). The homophone subset is the most challenging overall, indicating that phonetic similarity between $w_{p}$ and its replacement degrades the LLMs’ ability to distinguish them in context. This may also be influenced by orthographic similarity, as similar phonetics often imply similar orthography (e.g., stock/stork or dragon/brag gone in Example 3). Pun-word synonym substitutions (pun syn) also generally challenge the models, suggesting they may overlook the effect of the replaced word when it is semantically similar to a more fitting alternative. This suggests that LLMs might be actively “fixing” the pun in their reasoning process, thereby rectifying the alteration. The random and alt syn sets show slightly higher accuracy overall, further indicating that semantic or phonetic similarity to valid puns misleads predictions more than loosely related substitutions. These results suggest that LLMs fail to recognize contexts that should not trigger the pun effect in cases where the target sentence has a typical pun structure.
@@ Results.
Table 3 reports the average PPA over three runs, representing the mean number of words correctly identified in generated rationales for pun pairs on a 0–2 scale. GPT‑4o is the top performer, with $\sim$1.5 out of 2 correctly predicted words per pun across all three datasets. It is followed by Llama3.3 and DeepSeek‑R1, whose performance is noticeably lower. Mistral has the lowest scores, trailing the next-lowest model, Gemini2, by $\sim$33%. Among the 70B models, Llama3.3 performs best on average. We also observed no significant difference between the two prompts tested. Consistent with our earlier results, scores on the PunEval datasets are generally higher than on JOKER and NAP, where performance drops by approximately 20% and peaks at 1.5 out of 2. Furthermore, a PPA evaluation restricted to true positives (see Appendix Table 12) shows that most models, except Qwen and Mistral, can effectively justify their correct answers. In this specific evaluation, GPT-4o and DeepSeek-R1 emerge as the top performers by a wide margin. These findings motivate our investigation of the main error types in model reasoning, which is addressed in the next subsection.
@@ 6.3 Discussion
Our findings align with prior work showing that LLMs struggle with humor, particularly when it depends on nuanced context or roleplay Quan et al. (2025); Mirowski et al. (2024). These studies note that safety constraints (e.g., the “Harmless, Helpful, Honest” criteria) can cause LLMs to misinterpret relational context in potentially offensive situations. These safety mechanisms bias models toward a safer, “washed-out” form of humor, stripping away the surprise and edge on which many jokes depend. Puns, in contrast, are rich in surprise, subtlety, and timing; they are also deeply rooted in human emotions and experiences, qualities that current LLMs cannot fully tap into.
@@
Finally, the frequent pun pair errors point to weaknesses in assessing phonetic and orthographic similarity. These issues are likely tied to tokenization, which can mask morphological elements critical to wordplay, and to limited phonological modeling Liao and Shi (2025). The composition of pre-training corpora and data cleaning practices may also deprive models of the informal, creative language where puns are common. Because developers rarely disclose detailed preprocessing and corpus information, it is difficult to determine how these practices affect fluency with such language.
Improving an AI’s grasp of puns and humor in general will likely require deeper human-machine collaboration and greater social awareness. Prior work suggests that human alignment of LLMs should shift from a single global standard to a more granular, audience-aware approach Quan et al. (2025); Mirowski et al. (2024). This would help models to adapt to specific audiences and contexts, adjusting appropriateness and fairness criteria accordingly.
@@ Acknowledgments
We thank Felix Chadwick-Smith, Fayeja Farhana, Dhruvil Raval, Kayleigh Shier, and Hannah Trotman for their work as annotators on the manual error analysis of rationales. We also thank the entire Cardiff NLP group for their guidance on the project and for feedback on an early draft of the paper. Finally, we are grateful to Dr. Tristan Miller and co‑authors for kindly allowing us to use the JOKER dataset in our evaluation.
@@ C.3 RQ2: Pun Language Patterns
Table 10 compares the results obtained using each prompt for PunEval and PunnyPattern. Note that the best prompts (in bold) are the same as those that performed best in our RQ1 (Table 1). Fig. 11 visualizes the performance differences for each set of expressions containing language patterns. We interpret bias as being present in models with very high recall but low precision, indicating they tend to classify all instances containing the pattern as puns.
We observe a sharp imbalance of low precision and high recall for the never_die, used, and doctor patterns (Fig. 11), indicating that these patterns tend to mislead LLMs into identifying puns. In contrast, the tom and when patterns do not appear to consistently affect them. While all LLMs tend to produce many false positives with the used, never_die, and doctor patterns, results for the other patterns are more varied. For instance, GPT-4o, Qwen2.5, and Mistral3 are more resilient to the daughter pattern compared to the other models. Interestingly, examples containing the word “Tom” produce both false positives and negatives, suggesting that this pattern is unlikely to serve as a reliable shortcut for classifying puns. In contrast, sentences beginning with “When the” are classified more reliably by all models.
@@ C.5 RQ3
This targeted analysis reveals that most models can generate reasonably accurate rationales for the puns they successfully identify. Mistral and Qwen achieve lower scores in this setting, producing up to 40% incorrect rationales, while GPT‑4o, Gemini, and R1 produce correct predictions for the wrong reasons at least 20% of the time.